OpenStudy (anonymous):

find the equation of a line with slope of 2 and passes through (1, -3)

OpenStudy (texaschic101):

y = mx + b m = 2 (1,-3) x = 1 and y = -3 now sub your stuff in and solve for b -3 = 2(1) + b -3 = 2 + b -3 -2 = b -5 = b equation is : y = 2x - 5

OpenStudy (anonymous):

thank you so much :-)

OpenStudy (texaschic101):

no problem :)

OpenStudy (anonymous):

you rock :-) you are such a BIG help to me :-)